### v2

# RealitySummary: Exploring On-Demand Mixed Reality Text Summarization and Question Answering using Large Language Models

Abstract.
Large Language Models (LLMs) are gaining popularity as tools for reading and summarization aids. However, little is known about their potential benefits when integrated with mixed reality (MR) interfaces to support everyday reading assistants. We developed RealitySummary, an MR reading assistant that seamlessly integrates LLMs with always-on camera access, OCR-based text extraction, and augmented spatial and visual responses in MR interfaces. Developed iteratively, RealitySummary evolved across three versions, each shaped by user feedback and reflective analysis: 1) a preliminary user study to understand user perceptions (N=12), 2) an in-the-wild deployment to explore real-world usage (N=11), and 3) a diary study to capture insights from real-world work contexts (N=5). Our findings highlight the unique advantages of combining AI and MR, including an always-on implicit assistant, minimal context switching, and spatial affordances, demonstrating significant potential for future LLM-MR interfaces beyond traditional screen-based interactions.
Mixed Reality; Large Language Models; Augmented Reading; In-the-Wild Study; Diary Study
journalyear: 2024copyright: acmlicensedconference: arXiv; September, 2024; $c$s: Human-centered computing Mixed / augmented reality
Figure 1. We developed RealitySummary, an on-demand mixed reality reading assistant, over three iterations and conducted three different studies. Through this evolution, we highlight three key findings: 1) unique benefits of MR and LLMs, 2) emergent use cases of MR reading assistants, and 3) design implications for mixed reality reading assistants.
1. Related Work
We summarize prior work on reading augmentation and content summarization with a focus on interactive approaches.
1.1. Reading Augmentation
Document Enhancement
Human-computer interaction researchers have explored ways to augment physical documents with AR-enhanced content. Examples include HoloDoc (Li et al., 2019), Replicate and Resuse (Gupta et al., 2020), and PaperTrail (Rajaram and Nebeling, 2022), all of which laid the groundwork for augmenting physical paper by superimposing dynamic virtual content onto printed documents. Prior works have demonstrated these concepts on different platforms. For example, Pacer (Liao et al., 2010), MagPad (Xu et al., 2015), Tecahable Reality (Monteiro et al., 2023), and Dually Noted (Qian et al., 2022) use mobile AR for text augmentation, whereas Replicate and Resuse (Gupta et al., 2020) and Wiki-TUI (Wu et al., 2007) leverage head-mounted displays to overlay information. Other examples such as DigitalDesk (Wellner, 1991), DocuDesk (Everitt et al., 2008), Affinity Lens (Subramonyam et al., 2019), Matulic et al. (Matulic and Norrie, 2013; Matulic et al., 2013), and QOOK (Zhao et al., 2014) use projection mapping for document augmentation. At their core, such augmented reading research (Kirner et al., 2012; Billinghurst et al., 2001; Hidayat et al., 2020; Grasset et al., 2008) aims to improve cognitive load (Cheng, 2017), knowledge accumulation (Dünser et al., 2012; Zhang et al., 2019), and spatial visualization (Shelton and Hedley, 2004), given a long-standing consensus that readers prefer physical paper over screen-based reading  (Tashman and Edwards, 2011a; Sellen and Harper, 2002).
However, none of the existing systems have achieved an on-demand document enhancement. Instead, documents required pre-processing with content manually prepared before being embedded in AR. For example, HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), and Affinity Lens (Subramonyam et al., 2019) did not extract text directly from the given document; instead, content was prepared in advance and loaded based on attached fiducial markers. As a result, these systems are not fully adaptable and deployable to real-world scenarios. Some preliminary works have explored on-demand text analysis using machine learning, such as Dually Noted (Qian et al., 2022) analyzing document structure in real-time, Augmented Math (Chulpongsatorn et al., 2023) extracting graphs and math equations, and SOCRAR (Strecker et al., 2022) using OCR to extract key information. However, none have yet achieved comprehensive and general-purpose document enhancement.
Screen-Based Reading Support Tools
Outside of the context of mixed reality, a wide range of projects and systems have explored reading support tools for screen-based interfaces. Examples including LiquidText (Tashman and Edwards, 2011b), texSketch (Subramonyam et al., 2020), ScholarPhi (Head et al., 2021), Chameleon (Masson et al., 2020), and Marvista (Chen et al., 2022) developed on-screen reading support tools. Similar to our work, Marvista (Chen et al., 2022) helps users read online documents by providing reading aids including summaries, contextual prompt questions, and reading metrics for active reading. Other works studied active and close reading support like Metatation (Mehta et al., 2017) for closed reading, XLibris (Price et al., 1998) with free form annotation, Matulic and Norrie (Matulic and Norrie, 2012) to support navigation, and GatherReader (Hinckley et al., 2012) with information gathering. Other works also support literature review and reference search for science papers (Palani et al., 2023; Kang et al., 2023; Chang et al., 2023). Lastly, works like Explorable Explanations (Victor, 2011) and Potluck (Litt et al., 2022) try to enhance static documents with interactive content to improve reading comprehension through exploration. While these past works are limited to screen-based interfaces, this paper explores the augmented reading design for AR-based interfaces.
1.2. Content Summarization
Automatic Summarization
In the field of natural language processing, several works have employed automatic summarization to support readers (El-Kassas et al., 2021; Zhang et al., 2023). These approaches provide various summarization methods, including abstractive and extractive summaries generated from document photos (Benharrak et al., 2022), real-time summarization of user writing for quick text iterations (Dang et al., 2022), enhancing reading comprehension through automatic summarization (Chen et al., 2022), and responsive text summarization for adapting document summaries to screen sizes ranging from large displays to small watch faces (Leiva, 2018). With the recent advances in large language models like GPT-4, such automatic summarization tasks have become more accurate, robust, and accessible. For example, Goyal et al. (Goyal et al., 2022) showed that GPT-3 summaries were preferred by humans over those generated by fine-tuned models. Various types of summarization have been explored using GPT-3, such as opinion summarization (Bhaskar et al., 2022), news summarization (Goyal et al., 2022), and medical dialogue summarization (Chintagunta et al., 2021) while others have built custom frameworks for feeding long texts into GPT-3 to elicit more finely-controlled summarizations (Payne, 2022).
Interactive Summarization.
Researchers have also explored interactive summarization approaches based on context and users’ needs. For example, Hoeve et al. (ter Hoeve et al., 2020) studied Q&A interactions to assist users with a document-centric question-answering paradigm, finding that information-seeking was a common query. Semantic Reactor (Google, [n. d.]), developed by Google, enables users to ask questions about documents in natural language by fine-tuning BERT (Devlin et al., 2019) using the SQuAD dataset (Rajpurkar et al., 2018). VERSE (Vtyurina et al., 2019) is another Q&A system that allows users to ask specific types of questions about a document but is limited by its inability to summarize. Hoeve et al. (ter Hoeve et al., 2020) identified various question types asked by participants regarding documents, with factual/summarization questions being the most frequent. Some question types include document-related, factoid, mechanical (answerable by rule-based methods), factual, yes/no, navigational, and summary questions. Voice assistants like Cortana, Alexa, Siri, VERSE (Vtyurina et al., 2019), Firefox Voice (Cambre et al., 2021), and Pushpak (Holani et al., 2019) offer efficient query modalities, as speech input has been found to be significantly faster than manual input, especially for mixed reality devices (Ruan et al., 2016; Adhikary and Vertanen, 2021). Our goal is to build on top of this prior work, developing and exploring automatic and interactive summarization techniques for on-demand and real-time summaries embedded in documents.
The limitations of prior works such as Holodoc (Li et al., 2019), as explained before, prevented the study of long term and in-the-wild use of MR reading assistants. To our knowledge GazePointAR (Lee et al., 2024) is the closent work which explored long-term use of pronoun disambiguation in MR voice assistants. Still, it is unclear how can LLM powered MR assistants are used in everyday life to assist users in various types of reading and knowledge work tasks. We aim to fill this gap. In this paper, we contribute to the first demonstration of on-demand document enhancement by leveraging large language models. The goal of this paper is not only to demonstrate its potential but also to uncover its limitations and challenges from a user experience perspective through iterative implementation and multiple user evaluations.
2. Overview of Our Research Approach
Figure 2. Overview of our research approach spanning an evolution of our system across three different versions.
To uncover insights into on-demand reading assistants, we employed a three-phase iterative design process to investigate user interactions and experiences with RealitySummary. Our approach consisted of a usability study on the initial prototype (Study 1: RS-Doc), followed by an in-the-wild study on a second prototype (Study 2: RS-Wild), and concluded with an author-conducted diary study using a third prototype (Study 3: RS-Diary). Each phase was designed to build upon the insights gained from the previous round, allowing for continuous refinement of the design and functionality.
We chose this multi-method, iterative approach to capture a comprehensive understanding of user interactions across various contexts. The initial usability study provided controlled insights into the basic functionality and user experience of RealitySummary. The subsequent in-the-wild study allowed us to observe how users interacted with the improved, more reliable prototype in everyday environments, uncovering real-world usage patterns and challenges. Finally, the diary study offered deep, longitudinal insights into the integration of RealitySummary into daily life, allowing for nuanced observations of subtle interaction patterns and evolving user needs. This progression from controlled to naturalistic settings, coupled with iterative prototyping, allowed us to holistically explore the complex dynamics of AR-AI interactions while addressing emerging design considerations and ethical implications at each stage.
3. Study 0: Elicitation Study
Before implementing our system, we conducted a design elicitation study to explore potential designs for an AI-assisted augmented reading assistant. We held formative design workshops with six participants (4 males and 2 females) who had various design backgrounds as undergraduate and graduate HCI researchers. The study involved two tasks, each lasting approximately 15-20 minutes, with the entire session taking about 60 minutes.
3.1. Method
We asked participants to complete a four-part design task where they read and graphically annotated both a document of their own and a Wikipedia article [^1] Source: https://en.wikipedia.org/wiki/Russian_invasion_of_Ukraine. After reading, participants brainstormed ideas for AI-powered augmentations to improve readability. In the first design task, participants annotated a physical printout of their own article, sketching ideas for possible enhancements, and then repeated the task with a printed copy of the Wikipedia article. In the second task, participants were provided with digital copies of the two documents on a virtual whiteboard (Miro) and asked to design a second set of document enhancements, this time with assistance from ChatGPT.
Before the study began, we briefed participants, gave them a reference page of sample visualizations, and demonstrated prompts for ChatGPT that showcased its summarization and enhancement capabilities. Using a think-aloud protocol, the authors took detailed notes and collected all participant sketches. These designs were later collated on a Miro board, where four authors independently categorized and coded them, identifying 20 distinct types of document enhancements based on usage, purpose, popularity, and context. The authors then discussed and refined these categories, re-coding the examples to produce a final set of enhancement types  (Figure 3).
Figure 3. In our elicitation study, participants identified a variety of potential summary and augmentation types based on sketching and ChatGPT exercises using diverse text documents.
3.2. Results
We identified five high-level categories: summarize, compare, augment, extract, and navigate. The summarize category focused on condensing text for quick reference, ranging from broad document summaries to personalized summaries for different comprehension levels. Summaries were seen as useful reference tools, allowing users to quickly review content without losing focus. The compare category emphasized transforming textual and numerical data into visual formats like tables, graphs, and timelines, which could dynamically update based on user interactions, making comparisons more intuitive. The augment category explored ways to enhance document content with external or AI-generated data, such as adding maps and images or converting 2D figures into 3D models. Advanced ideas included visualizing the importance of academic citations or integrating external information to provide richer context. The extract category focused on enabling users to pull out and store key elements, such as keywords, figures, or quotes, for future reference. Lastly, the navigate category proposed various methods to improve document navigation, such as progress indicators, accessible tables of contents, and using AR to make documents searchable. These insights, alongside considerations of the potential and challenges of AI-driven tools, informed the development of our system.
4. Study 1 RS-Doc: RealitySummary First Prototype and Usability Study
We implemented the first version of our system on a Microsoft Hololens 2 and conducted a preliminary user evaluation. In this section, we present initial insights from using on-demand, mixed reality document summarization and discuss the associated limitations and design implications.
4.1. Implementation
Figure 4. RS-Doc Hololens 2 architecture.
Figure 5. Examples from our RS-Doc Hololens implementation.
The system design for the HoloLens version involves capturing documents using the HoloLens 2’s internal camera and extracting text with Google Cloud OCR. The extracted text is processed using GPT-4 via OpenAI’s API to identify key information, while spaCy’s named entity recognition (NER) module assists in identifying important entities for further web searches. The system employs Vuforia for markerless image tracking and the Mixed Reality Toolkit 3 (MRTK3) SDK for content rendering. Users interact with the system through voice commands processed via Google’s speech-to-text API, allowing them to modify or query visualizations. The system captures images periodically for OCR, with an option for manual activation, and tracks documents by generating dynamic image targets, ensuring accurate document tracking and augmented content rendering.
Informed by the elicitation study, the system supports several document enhancements aimed at improving the reading experience. These enhancements include text summarization, achieved by sending the current page’s text to GPT-4 for a concise summary, limited to 150 words for easier digestion. Comparison tables are generated on demand based on the document parsing, allowing for quick comparisons within the document. Timelines of events are created by prompting GPT-4 to generate a sequence of summarized events. A keyword list is extracted using GPT-4 and spaCy, with spaCy handling named entities for web searches and keyword highlighting. Summary highlighting allows users to obtain brief summaries or additional information on specific phrases, sentences, or paragraphs, which are then visually highlighted. Information cards provide further details on keywords, people, places, or images, leveraging Google Search, Image APIs, and Mapbox for generating maps. All these features are integrated into the HoloLens environment, offering real-time, interactive content within the mixed reality experience.
4.2. Method
The usability study involved 12 participants (P1-P12, 6 male and 6 female, aged 19-35) recruited via word-of-mouth and compensated \$15 for their one-hour participation. The study aimed to evaluate the system’s usability, explore participants’ preferences for different document enhancement strategies, and identify limitations for future iterations with real-world use in mind. Conducted in a lab environment with Microsoft HoloLens 2, participants were first introduced to AR and the HoloLens 2, followed by an overview of the system and its document enhancements. They completed three reading tasks using different document formats (physical, digital, and a book) and utilized the system’s enhancements as they preferred. The participants’ digital documents included books, blog posts, course assignments, Wikipedia articles, magazine articles, and news articles viewed on various devices.
Throughout the study, participants were encouraged to think aloud, and their experiences were documented. Following the tasks, semi-structured interviews and a Likert scale questionnaire, including a System Usability Scale (SUS), were conducted. Thematic analysis was performed on the transcriptions of the think-aloud sessions and interviews, resulting in the identification and refinement of key themes.
4.3. Results
Overall, participants responded positively to RealitySummary, with most participants agreeing that the prototype was easy to use (M=4.3/5, SD=0.98), easy to understand (M=4.5/5, SD=0.67), and well-integrated with the task (M=4.3/5, SD=0.77) with 5-point Likert-scale questionnaire. Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with comprehension (M=4.6/5, SD=0.514), were relevant to the document content (M=4.7/5, SD=0.65), improved comprehension speed (M=4.4/5, SD=0.79), and augmented their reading experience (M=4.7/5, SD=0.65). Finally, the System Usability Scale (SUS) returned an overall composite score of 71, suggesting a reasonable level of usability for an early-stage prototype.
4.4. Insights
Our analysis identified several themes related to participants’ experience in comprehending and navigating documents. We also explore participants’ comparisons of the visual and text-based summarization techniques as well as preferences for various augmentation techniques.
Enable Multi-Angle Content Understanding for the Current Page
RealitySummary’s document enhancements can significantly improve readers’ comprehension. Several participants (P7, P8, P11, P12) highlighted that features such as summaries, timelines, keywords, images, and people cards provided a more comprehensive understanding of the document, allowing them to engage with the content from multiple angles and perspectives. All participants mentioned that RealitySummary enabled them to grasp content more quickly than with their usual methods, such as visually scanning a physical book or manually searching for timelines on a phone or PC. For instance, after reviewing the first magazine, P12 remarked, I haven’t read this magazine thoroughly, but it seems these features can significantly help me understand what it’s about in less time.” Interestingly, several participants appreciated that the summaries and keywords were drawn only from the currently visible page rather than from the entire document. P5 reported, I like the page-by-page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document.” Similarly, P6 emphasized, “I like that it (the system) gives you keywords for each page and not for the whole document.
Facilitate Efficient, Top-Down Navigation
Both structured overviews and in-context highlights provide useful entry points that make it easier to navigate documents from a top-down perspective, even within individual pages. Broadly, participants spoke positively about how RealitySummary helped them navigate documents. Four participants (P3, P4, P6, P8) specifically mentioned using structured navigation elements like timelines to better understand the structure and flow of the text, with P4 noting that the timeline helped me organize and navigate my thoughts very quickly while reading the text.” In fact, ten participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining the high-level ideas and terms, then transitioning into reading the lower-level content as needed. P8 highlighted, “[the enhancements] give you a foundation before you actually jump inside [the document].” This approach enabled participants to prioritize their reading, allowing them to “focus on things that you’re more unsure about” (P1).
Support Quick Understanding with Visual Aids
Both visual and text-based enhancements offer distinct benefits to readers—visual aids provide quick, high-level overviews, while text summaries capture more complex details. Some participants preferred text-based features for gaining a deeper understanding of specific topics, while others found that visual elements improved their comprehension speed, engagement, and retention. All 12 participants rated the visual summarization techniques as helpful, with eight specifically highlighting the usefulness of images and bio cards. Several participants (P1, P4, P6, P7, P9, P10, P11) noted that visual augmentations allowed them to grasp the document’s content more quickly compared to reading text alone, with P11 stating, “Seeing the visuals helped me grasp what’s happening in the document right away.” Participants (P7, P8) also emphasized that visuals enhanced their engagement with the material, with P7 commenting, I’m not a big fan of reading, so having visuals helps me engage with the content.” Additionally, participants highlighted the potential of visuals to improve memorability, with P8 noting that for historical documents, people cards helped put a face to a name.” Similarly, P7 observed that while text-based summaries aided understanding, “graphical elements help me remember it for a longer time.”
4.5. Limitations
Slow and Unreliable Document Tracking
Document tracking enables unique features like keyword highlighting and around-document placement. However, the primary limitation of the current prototype is the instability of its document tracking. Overall, participants reported dissatisfaction with the system’s slow, inaccurate, and unreliable tracking capabilities. Unlike the OCR process, document tracking involves a heavier and slower workflow, requiring rectangle detection, image extraction, and image upload for Vuforia processing. Even with a perfect image target, tracking performance varies based on several factors, such as lighting conditions, hand occlusion, and viewing angles. Furthermore, while the tracking performed better for visually rich documents, it struggled with text-heavy documents, as it does not have any distance features, making consistent tracking difficult over time. Even when tracking functioned properly, participants frequently experienced disruptive flickering, which negatively affected the user experience. Additionally, the fixed placement of augmented information on documents posed a challenge for long-term use, as participants expressed the need for more flexibility in adjusting the location of the augmentations.
Layouts and Visual Clutter
Several participants (P7, P10, P11) noted that around-document augmentation helped them focus on their reading but emphasized that the number of displayed cards and their placements should be adjusted based on the user’s needs and context. Displaying too many cards at once can lead to visual clutter and divert attention from the document itself. We observed that participants might become overly reliant on the generated content rather than engaging directly with the reading material. Whether positive or negative, 11 out of 12 participants reported that they could understand the document without actually reading it, relying entirely on the contextual information provided by RealitySummary. P9 even stated, “I didn’t even read the document, but I understand what it’s about.” Currently, the system automatically generates various types of cards when relevant, but allowing users to reduce the number of visible cards based on their preferences could help manage visual clutter and maintain focus.
Fixed Prompt and Augmentation Types
Another key issue was the inability of participants to query the generated summaries. The fixed prompt approach used for generating the different types of cards limited the system’s flexibility in addressing broader use cases and contexts. Several participants (P1, P3, P5) expressed frustration with this constraint, as they had specific questions they wanted to ask but were restricted to the preset summary types provided. While these different types of augmentations offered different benefits, those advantages were only realized when the generated content was relevant and accurate. Participants also noted that specific aids, like timelines, comparison tables, and keyword lists, were not always helpful or contextually appropriate. On the other hand, participants generally appreciated the broader usefulness of the general summaries and visual information cards, which were more adaptable to varying contexts.
Applicability Beyond Controlled Environments
Finally, the controlled nature of the study limits its ability to assess the broader applicability and real-world uses. While the initial findings are encouraging, they do not fully reflect how the system might perform in diverse, unstructured environments. As a result, transitioning to in-the-wild studies will be essential to understanding how document augmentations work in everyday scenarios and to gain insights from prolonged, regular use in real-world settings.
In conclusion, a more refined version of the system, featuring flexible augmentation placement and query-able summaries, is essential for better evaluating its effectiveness in real-world scenarios.
5. Study 2 RS-WILD: In-the-Wild Study with the Second Prototype
Figure 6. Examples from our in-the-wild study.
To explore real-world use cases, we conducted an in-the-wild study with eleven participants. Unlike the previous usability study, which focused on standardized reading tasks in a controlled lab setting, this study aimed to examine the versatility and applicability of our system in uncontrolled environments and with diverse reading materials. The goal was to uncover unique advantages and potential use cases for using mixed reality and AI-generated assistants in everyday settings. For this study, the system was deployed on an Apple Vision Pro, offering considerably higher rendering quality and a broader field of view compared to the Hololens 2. This contributes to better readability, particularly for text summarization, by using a less distractive transparent background. Based on the user feedback from the usability study, the user interface was simplified with a single text box and an image slot, which could accommodate any response type given the user’s custom query. This feature would better suit diverse applications and minimize the intrusiveness of additional content in daily use.
5.1. Implementation
The system design for the Vision Pro follows a similar approach to the Hololens but with key adjustments. Since the Vision Pro does not yet allow direct access to the camera view, we implemented a workaround for text extraction. First, we screencast the Vision Pro to an external iPad, which is then connected to a MacBook Air to stream the feed via QuickTime Player’s mirroring feature. A web-based interface on the Mac captures the live stream from the Vision Pro. Using the captured feed, we employed Google OCR to extract text, which was then processed by GPT-4 via the OpenAI API. The Vision Pro application, developed in Swift, featured a single movable card with an image/map slot. After querying the LLM, data was transmitted from the web-based interfaces to Vision Pro via Firebase’s real-time database.
Additionally, a separate web interface with speech recognition was run on the user’s smartphone, allowing them to query the system via voice (using Web Speech Recognition) and ask follow-up questions. This enabled users to request specific summary types or submit open-ended queries about the text they were reading. The web-based interfaces, built with React.js and hosted on Glitch, were accessed by participants through their phones.
5.2. Method
We recruited 11 participants (W1-W11, 7 male and 4 female; ages 19-32 years) from the local community. They were compensated \$25 for participating in a 1.5-hour study, which included a one-hour experiment and a 30-minute interview. The study took place in various locations at a local university, such as a library, cafeteria, store, hallway, classroom, and multiple labs.
Before the experiment, we collected participants’ demographic information and reading habits, which were input into the system as prior knowledge for ChatGPT. Participants familiarized themselves with the Apple Vision Pro by completing the eye-tracking calibration setup. Unlike structured tasks or prepared reading materials, participants were encouraged to use the system with any text they encountered during the study, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment. Throughout the one-hour experiment, participants were encouraged to think aloud about their experiences. After the experiment, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the one used in the earlier usability study. Both the interview and questionnaire took approximately 30 minutes. As in the previous usability study, we analyzed participants’ questionnaire responses and reflections to identify high-level trends and opportunities for augmented reading tools.
5.3. Results
We gathered feedback from participants on various aspects, including the reading experience (immersiveness, understanding, distraction), perceptions of the system (content and query comprehension, trustworthiness, and privacy concerns), and overall usability (content placement and comfort using the MR headset). In general, participants reported feeling more immersed and interested while reading (M=3.5/5, SD=1.04), and they perceived the system as effectively interpreting their queries (M=3.3/5, SD=1.00). Besides, their responses on the content placement within the system were also overall positive (M=3.4/5, SD=1.29).
Figure 7. In-the-wild study questionnaire responses.
Participants had varied opinions on whether the system became distracting throughout the reading process (M=2.6/5, SD=1.37) and whether they tended to trust the accuracy and reliability of contents provided by the system (M=2.8/5, SD=1.25). Privacy concerns were also noted by the majority (6/11), and discomfort using the headset was prevalent among the respondents (8/11). These findings are detailed in Figure 7.
Figure 8. Snapshots captured from the Vision Pro during our in-the-wild study.
5.4. Insights
Observations from the in-the-wild deployment highlighted a wider range of use cases for augmented reading as well as some of the practical trade-offs associated with real-world use.
Types of Reading Content
The in-the-wild study demonstrated our system’s broad utility, revealing diverse exploratory use cases across various reading materials. Participants engaged with a variety of books, spanning textbooks, encyclopedias, fiction, and non-fiction. The system was also employed for different types of documents, including academic papers, handouts, and class assignments, as well as digital content accessed through smartphones and laptops, like social media posts, online news articles, and e-commerce websites. Beyond conventional reading materials, participants also used our system with everyday objects and texts in the environment, such as restaurant menus, printer instructions, product nutrition labels, furniture assembly guides, posters, and signage on trash bins and advertisements. Notably, some participants even applied our system to their handwritten notes from lectures.
Benefits of Always-On AI Assistant
The exploration of such diverse reading materials can be attributed to the always-on capability of the mixed reality reading assistant, as opposed to screen-based interfaces. A key benefit of this feature is its support for implicit inputs, which dramatically lowers the barrier to asking questions of LLMs. This automatic, continuous content capture broadens the range of applications beyond what participants initially envisioned. For instance, W4 mentioned they had never thought to type a question into ChatGPT when looking at a product label. Even for common reading tasks, they appreciated that the summary was automatically generated and updated in the background as they flipped through pages in a library. Overall, the always-on, implicit assistant made it effortless to integrate AI into daily tasks, expanding potential use cases and drastically reducing the explicit effort required from users.
Unique Affordances of Spatial and Tangible Interactions
Participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration. In a library setting, for example, participants could naturally navigate through bookshelves (spatial exploration), physically pick up a book (tangible interaction), and browse through its pages to automatically receive summaries (always-on assistance). W5 and W10 even leveraged our system to provide summaries or recommendations for entire collections of books on a bookshelf, not just content from a single page. Furthermore, W1 demonstrated the capability to synthesize information across multiple books on a table, requesting summaries or connections between various books and pages. Such observations highlight how mixed reality interfaces enable spatial and tangible interactions that go beyond the limitations of traditional screen-based interfaces, which can only capture and summarize visible content.
Unique Ways of Using AI-Generated Content
We observed several unique and unexpected use cases that took advantage of the LLM’s capacity for free-form questioning and answering. For instance, when reading a physics textbook, W8 asked the system to generate ten quizzes to help him understand quantum mechanics, transforming the textbook into a more active and personalized learning tool. Other common usages included translations (W4, W5, W6), learning aids (W4, W8), and personalized content recommendations (W5, W10). For example, W8 applied the system to his handwritten calculus notes to find answers to particular questions. Beyond traditional reading applications, the AI’s capabilities supported practical decision-making; W4 compared nutritional information of products on store shelves and menu items to determine healthier options, while W5 used it for guidance on waste disposal and printer operations. The LLM’s ability to process a broad spectrum of inquiries clearly expanded the system’s applicability and utility.
Proactive Summaries vs On-Demand Question Answering
Our system offers both 1) proactive assistance through automatic summary generation and 2) on-demand assistance where users manually ask questions. Participants shared their insights on these two functionalities by comparing the benefits and limitations. Proactive summaries were appreciated for providing a useful initial overview of unfamiliar topics. Preferences for proactive summaries varied with reading habits. For example, infrequent readers like W6 and W10 found these proactive summaries helpful in deciding whether to read the content or understand it without reading the full text. On the other hand, W10, who is a more frequent reader, found the constant summary updates unnecessary. In contrast, on-demand question answering allowed for a more focused engagement on specific topics. For instance, W1 enjoyed the “rapid-fire” questioning with a physics textbook, asking more questions than with traditional ChatGPT saying ”You have to describe it properly you have to specifically say what you’re asking about what you’re reading, so that kind of all of this abstracted in this system is like frees up your mind to just focus on the curiosity part Like rapid-fire questions and get rapid-fire answers”. W2 valued how questioning was woven into reading, preserving thought flow without the disruption of toggling between reading and ChatGPT interfaces. Overall, participants favored the on-demand feature due to its applicability to diverse situations, yet many appreciated having both options for different purposes: proactive features for implicit skimming and on-demand for addressing more specific questions.
Potential Risks of AI-Driven Augmented Reading
Participants also expressed several concerns about potential risks, notably privacy issues due to the possibility of constant surveillance. Despite this concern, most participants were willing to continue using the system, if they could control when it was active. For example, W1 expressed the desire to capture all text throughout her entire life, allowing for an extensive searchable archive of everything she has encountered. Additionally, participants also mentioned the risk of decreased motivation and skill development. For example, W6 mentioned that relying too heavily on the system for translations could reduce the motivation to learn new languages. Similarly, W11 worried that immediate access to answers for every question might contribute to a broader intellectual complacency. Overall, participants wanted the flexibility to choose when and how the system would be used to mitigate these risks.
Trust of AI-Generated Content
Many participants approached the system’s responses with caution, often verifying the answers with the reading material. They highlighted the necessity of consulting additional sources for in-depth or technical readings when depending on the system. There was a general agreement among participants about their reluctance to fully trust an always-on reading assistant in their daily activities. Conversely, we noted some intriguing discussions regarding how participants perceive our system, either as a mere tool or more like a personal companion. While most participants engaged with the system as a conventional chatbot, some regarded it more as a personal companion than a mere tool. Notably, W3 and W5 described the system as a friend for discussing and inquiring about their readings. The system’s ability to provide always-on, immediate responses facilitates this personification.
5.5. Limitations
Overall, the system design of the second prototype enhanced its applicability and stability for real-world use. Compared to the first prototype on the Hololens 2, the Vision Pro’s superior rendering quality and broader field of view provided a better user experience. Additionally, the simple system design greatly improved usability despite the absence of document tracking. While the second prototype showed great potential, we also identified several feature-level limitations, along with inherent study-related constraints.
Lack of History and Recall Support
While the minimalist design effectively reduces visual clutter and conserves screen space, it also limits the system’s overall utility. Although users appreciated the clean and simple interface, many expressed frustration with the lack of options to save or revisit previously generated responses. Several participants suggested that a history feature or the ability to store responses for future reference would substantially enhance the user experience.
Challenges with Screen Viewing for Extended Use
Many participants raised concerns about the difficulty of clearly viewing phone and computer screens through the Vision Pro without dramatically enlarging the text. While this worked for short-term tasks like quickly reviewing an assignment, participants found it frustrating for longer viewing periods, particularly when using laptops. Since much of their work is done on computer screens, this limitation poses a challenge for assessing the system’s effectiveness in supporting realistic, work-related tasks.
Need for a Longer-Term Study
Our in-the-wild study was inherently limited by its short-term, public use — with each session lasting only one hour per participant. As a result, participants did not have the opportunity to fully engage with the system in their personal workflows. For instance, due to privacy concerns related to screen recording, participants rarely used our system for more personalized content, such as emails, messages, or calendars. Furthermore, since participants were encouraged to explore various scenarios, they tended to experiment broadly rather than focusing on any specific task in depth. While the in-the-wild study provided valuable insights into real-world use cases, it only captured short-term interactions in public spaces. There is still much to learn about the sustained, regular use of reading assistants in daily life, particularly in how users interact with information over time, revisit previous queries, or support ongoing tasks.
6. Study 3 RS-Diary: Diary Study with the Third Prototype
To explore the long-term use cases, we developed a third iteration, incorporating feedback from the in-the-wild study. This version introduced a multiple responses feature and added the ability to view or mirror the content on an external device, addressing the readability issues for the long-term use of the Vision Pro.
Figure 9. Snapshots captured from the Vision Pro during our diary study.
6.1. Implementation
The final version of RealitySummary builds on the earlier Vision Pro implementation with two key improvements: 1) Users can now add multiple draggable cards on-demand by either making a new query or clicking an “add” button. User can spatially anchor these cards anywhere in their immediate environment and can also re-select cards to update them with new prompts or text from the scene. These cards are stored as JSON objects in a real-time database. 2) The system allows users to mirror their device screen within the Vision Pro for high-resolution viewing. This functionality — implemented through a WebRTC video call or via Mac Virtual Display on the Vision Pro — lets users view their external devices at full resolution in the MR environment and also allows RealitySummary to more accurately capture and process on-screen content.
6.2. Method
Five authors (A1-A5) used the system over the course of 3 weeks, integrating it into their day-to-day activities and accumulating 30 total hours of use. Due to physical fatigue and battery limitations of the Vision Pro, each session lasted less than 2 hours. The authors independently documented their use cases, insights, frustrations, thoughts, and reflections while using RealitySummary, which we later thematically coded. During the study, each participating author recorded their observations via a phone microphone, following a think-aloud process. Additionally, the first author conducted follow-up interviews to gain deeper insights into their experiences.
6.3. Results
Throughout the diary study, participants used the system in various settings, including labs, kitchens, outdoor areas, and offices, though the majority of usage took place at desks in work environments. They engaged with a wide range of reading materials, such as academic papers, news articles, code, emails, calendars, Google Scholar, classroom assignments, lecture slides, textbooks, and YouTube videos. Participants quickly adapted the system to fit a variety of personal, academic, and technical tasks. For example, A4 used the assistant not only to summarize email content but also to gather higher-level insights, such as identifying how many emails were sent by a particular sender. A2 utilized the system to support programming assignments, where they used the system to break down complex coding problems into manageable steps and provided detailed explanations of specific code snippets. A2 also leveraged the system for learning aids, such as generating to-do lists and quiz questions based on problem descriptions. The ability to spatially organize tasks and concepts allowed him to effectively manage his workload, maintain focus, and dynamically track his progress. Additionally, in his language learning efforts, A2 used the assistant to generate contextual vocabulary lists from Japanese reading materials, which helped him focus on the most relevant and frequently used words. A5 employed it to extract a step-by-step sequence of keyboard commands from an article on rebooting a MacBook. Overall, the diary study highlights the system’s flexibility in supporting diverse tasks, from email management and coding to technical troubleshooting, language learning, and academic research.
6.4. Insights
From these experiences, participants shared various insights, frustrations, and potential future improvements for the current prototype. Our diary study also provided insight into the use of an always-on reading assistant in more realistic work-oriented settings.
Quick and Simple Summary Responses Meet Many Immediate Needs
Participants valued the system’s ability to deliver quick, concise responses, which were often sufficient for immediate clarifications. A2, for instance, appreciated how the system provided brief answers when they needed a rapid comparison of philosophical concepts, like distinguishing between act and rule utilitarianism. In many cases, just a high-level summary captured through manual input was enough to satisfy their needs. A4 found this method to be a useful shortcut, as pressing a button was faster and less disruptive than speaking and often provided exactly the information they were seeking. These simple, proactive summaries were particularly beneficial for tasks that required quick comprehension and minimal interruption.
Minimal Context Switching Enhances Focus and Flow
Participants appreciated the integration of LLM into the MR interfaces, in part, because it helped them minimize context switching, allowing them to maintain greater focus. The system’s automatic text extraction feature enabled participants to stay in a flow state. For example, A2 noted that the interface allowed them to concentrate on their work without the mental disruption of switching between tasks or manually inputting context, even when using a computer screen. This seamless workflow kept participants engaged, enhancing productivity and reducing cognitive load. A1 highlighted how the system functioned like a “teacher” by offering real-time assistance without requiring them to step away from their tasks. Participants also remarked that on-demand support boosted their efficiency during complex tasks like coding or writing, where focus is critical.
Using Generated Cards as Peripheral References
Participants highlighted the advantages of peripheral output modalities. For instance, A2 used the system as a reading aid while learning Japanese. Instead of frequently opening a dictionary, which disrupted the reading flow, they generated vocabulary flashcards around the document. This provided the needed language support without interrupting the reading and learning process. A4 also noted that they would often cluster and group these generated cards thematically to the sides of the display for quick reference and then delete them once they were no longer needed. Similarly, A1 used the system to “pin” generated code versions and key concepts around their workspace, allowing for easy access without breaking focus and enabling them to quickly revisit important points as needed.
Spatial Organization Improves Recall and Multitasking
A key finding from the study was how participants leveraged the system’s spatial organization features, similar to arranging post-it notes on a board. Both A1 and A2 utilized this capability to group related concepts together. For example, A2 clustered cards on CPU architecture, placing related topics like buses and registers nearby, much like organizing post-its by theme. This spatial arrangement enabled quick access and improved recall, as users could visually track where specific information was placed. Similarly, A1 pinned cards documenting different problem-solving methods across their workspace, ensuring easy access when needed. This post-it-like organization supported memory retention and task management by allowing users to arrange information in a way that aligned with their thought processes.
6.5. Limitations
Challenges Controlling What to Capture
Currently, the system captures the user’s entire field of view, but participants expressed a desire for more control over what is captured. For example, A1 struggled to get the system to ignore personal emails displayed next to the article he was reading, leading to summaries that inadvertently included private information. Similarly, A2 wanted to focus on a specific code snippet to obtain more accurate results while debugging. Additionally, the lack of visual feedback left users uncertain about what OCR content or context was being processed in each request. To address these issues, the system could offer enhanced visual feedback and allow users to control specific areas of the screen, such as through gaze interaction, which would help mitigate privacy concerns and prevent the inclusion of irrelevant or personal content.
Cumbersome Re-Prompting Process for Common and Repeated Tasks
Participants frequently found themselves re-prompting similar queries for routine tasks, such as requesting bullet point summaries or filtering by specific criteria. A potential improvement could be the addition of dynamic buttons that save commonly used prompt types, allowing users to easily reuse them later. Additionally, participants expressed frustration with testing out lots of different prompts to generate a complex output, such as aggregating the Google Scholar page into a timeline. Saving or sharing the same specific instructions for repeated tasks might mitigate these problems.
Managing Growing Numbers of Summaries
In the first prototype, participants found it distracting to have multiple summary cards anchored to the image in their workspace. However, they did not find this frustrating when they had control over the placement and could easily delete the cards. Additionally, participants noted that the system could sometimes generate an overwhelming number of cards. For instance, A2 mentioned that while analyzing a poem, the system created a card for each sentence and its meaning, making it difficult to manage information overload. This feedback suggests that there is potential value in exploring new interactive and semi-automated methods for efficiently sorting, snapping, organizing, filtering, and removing these cards. Such improvements could enhance the usability of the system and help users better manage the content generated.
7. Lessons Learned from Three Studies
Drawing on results from across our prototyping and evaluations, we close by showcasing the unique benefits of combining MR and AI tools, enumerating potential use cases, and highlighting opportunities for future research.
7.1. Unique Benefits of Combining MR and AI
We identified two unique benefits of combining MR with AI: 1) implicit input and 2) peripheral output.
Always-On Implicit Input
The always-on camera in MR enables implicit input, distinguishing it from the explicit input required in desktop interfaces, such as opening an app, typing queries, or copy-pasting contextual text. Participants in in-the-wild and diary studies appreciated the system’s ability to capture information without requiring explicit effort. This always-on input allows for hands-free exploration in both spatial and temporal contexts, such as scanning entire bookshelves as a query or utilizing the history of flipped pages. Implicit input reduces cognitive load and minimizes context switching, as users are not required to actively specify what information should be used for the LLM’s context.
Peripheral and Background Output
We found that MR’s spatial output allows information to be presented in a peripheral, background manner. In contrast, other output modalities, such as audio or screen-based displays, often require the user’s foreground attention to consume generated content, disrupting focus and interrupting workflow. With MR and peripheral outputs, however, AI-generated content becomes less intrusive, as it can be easily ignored if not needed, simply by choosing not to look at it. For example, participants in the diary study often appreciated this feature for learning aids, such as understanding textbooks or accessing always-on vocabulary references when learning a foreign language. This peripheral output is a unique aspect of the integration of MR and AI. We believe this capability would promote more calm human-AI interactions in the future, reducing disruptions to focused attention by combining implicit input with unobtrusive background support.
7.2. Emergent Use Cases for Mixed Reality Text Augmentation
Both the in-the-wild and diary studies revealed a variety of interesting emergent use cases. To provide a more comprehensive view of how users engage with these systems, we have categorized these into seven groups:
Summarizing Content
The primary use case was to summarize content. This was the most commonly observed behavior across various contexts. Participants summarized a wide range of texts, including research papers, blog posts, news articles, books, and instructional guides.
Extracting Information
Participants also utilized the system to extract key information, which can be translated into a more consumable format. This aligns with design elicitation results, which suggest tables of contents, keyword lists, timelines, and comparison tables. For example, one participant extracted vocabulary flashcards to assist with foreign language reading, while another participant created a table of keyboard shortcut commands from a troubleshooting guide to use as a reference. In these examples, the consumable format, such as tables or vocabulary lists, serves as helpful references for their tasks.
Querying External Sources
Participants frequently used extracted text as a query to retrieve additional information from external sources. Echoing the information types identified in our elicitation study, participants queried specific names or keywords from their reading materials. For instance, one participant retrieved details about a player, team, and standings from a sports-related article, while another fact-checked claims on-demand when reading news articles.
Asking for Suggestions
Beyond the above use cases identified in the elicitation study, participants frequently asked for recommendations or suggestions based on the current context. For instance, users asked for book recommendations when looking at a bookshelf. Beyond the reading context, one participant asked for advice on which trash bin can be used based on disposal instructions, while others inquired about healthier menu options or products. They also asked for recommendations from external sources. For instance, participants asked for suggestions for relevant research papers when reading academic articles.
Aggregating Information
An interesting use case involved using the system to aggregate high-level information. Unlike simple information extraction, aggregation often involves operations like filtering and counting. For example, some participants used the system to count emails based on specific filters, while others filtered and counted first-author papers from a researcher’s Google Scholar page. This allowed users to perform quick, approximate queries that would normally require more complex searches.
Transforming Content
Participants also used the system to transform their reading content into interactive, actionable outputs. For instance, some participants applied our system to generate quizzes from class notes or textbooks, while others decomposed high-level assignments into step-by-step to-do lists. One participant asked the system to create reminder cards that remained spatially persistent. Participants also use our system to translate from one language to another. Unlike information extraction, these tasks involved converting content into actionable, interactive outputs.
Generating New Content
Finally, participants generated new content based on existing text. In one example, a participant requested multiple variations of a written paragraph, which they then organized spatially. Others used the system to create code based on instructions or to debug existing code snippets.
These emergent use cases highlight novel ways of interacting with AI-generated information, allowing users to engage with individual snippets and convert them into dynamic outputs. They also illustrate how participants adapted the assistant to handle different types of content, which suggests a broad space of possible designs for such systems.
7.3. Future Opportunities and Design Implications
As both MR and LLM technologies continue to evolve rapidly, considerable opportunities exist for tools that more richly integrate them. Based on our three studies, we identify several key opportunities for future research and design.
Enabling Greater Control Over Capture
While the system’s implicit input offers significant benefits, it also raises privacy concerns due to the automatic capturing without the user’s constant awareness. To address this, future designs should give users control over what data is captured. As noted in the diary study, participants expressed a desire for more transparency and control over the captured content. One approach could involve providing visual feedback, allowing users to see and select the specific areas or information being captured. Future systems should balance the benefits of implicit input with privacy considerations by offering user control and visual feedback.
Improving Organization and Filtering of Interactive Cards
Participants expressed the need for more efficient ways to manage the growing number of spatially organized content cards. While some diary study participants manually arranged these cards around documents, others suggested more advanced organizational tools. Features like sorting, snapping, filtering, or automatically organizing cards could simplify workflows and reduce clutter. Such capabilities would be particularly beneficial for long-term tasks, where the accumulation of content can become overwhelming. Integrating these organizational tools would enhance usability, supporting users in managing dynamic and task-specific information over time, which enables better workspace management for daily tasks.
Balancing Between Reusable Prompts and On-Demand Custom Queries
Our studies highlighted the importance of accommodating both reusable prompts and custom queries. Each has its advantages: prepared prompts can generate complex outputs such as timelines or comparison tables, while custom queries offer users flexibility in generating content. The diary study participants expressed a desire for dynamic interfaces that could quickly reuse and adapt prompts to different contexts. For example, some suggested having customizable buttons that automatically generate summaries or lists in a consistent style. Others mentioned the need for a more efficient way to reuse queries across tasks, such as repeatedly generating lists or comparisons for class topics. This feedback highlights the importance of designing adaptive interfaces that allow users to seamlessly reuse and modify their prompts, improving efficiency and user satisfaction.
Facilitating Seamless Transitions Between Screen and Space
Participants also valued the ability to drag content cards from the system into their computer, such as pasting generated answers directly into coding tasks. Some participants also envisioned this feature being useful for writing tasks, allowing for a smooth transfer of generated content. Future systems should explore leveraging cross-device interactions between MR and other devices to facilitate seamless transitions, enhancing the workflow between screen-based and spatial interactions.
Expanding Beyond Text Input to Capture Images and Everyday Objects
Participants expressed a strong desire for the system to recognize not only text but also images. For example, the in-the-wild study revealed the limitations of inquiring about a diagram that the system could not recognize. Users also anticipated broader applications beyond text recognition, such as generating recipes based on ingredients in the fridge or providing information on real-world objects. Enhancing the system with multi-modal LLM capabilities could address these needs, expanding its utility beyond text-based interactions.
Personalization Through Always-On History
Lastly, participants showed interest in an “always-on” history feature that would allow them to recall previous interactions with the system, creating a more personalized experience. They speculated that such a system could eventually serve as a personalized assistant, continuously capturing and understanding context through an always-on camera. Although current hardware limitations, such as battery life and physical fatigue, prevent continuous long-term use, we believe future MR interfaces will seamlessly integrate with AI to provide both context-aware and personalized assistance based on users’ daily activity histories.
8. Conclusion
RealitySummary is the first on-demand MR reading assistant that combines OCR and LLM technologies to generate spatially-situated on-demand reading aids for both physical and digital documents. Through our empirical studies, we aimed to highlight new possibilities for on-demand, AI-powered reading support. This paper presented an iterative development of the system, along with key findings and lessons learned from three studies. We hope our work will inspire the HCI community to further explore the potential of integrating MR and AI in future applications.